A Bibliography of the Antikythera Mechanism

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

19 October 2024 Version 0.12

Title word cross-reference

2000 [36]. $8.50 [14].

-Year-Old [36].

000-year-old-computer [43].

10 [64].

2013 [113]. 2014 [64, 66]. 21st [113].

80 [5, 3, 4]. 85181 [30].

978 [30]. 978-0-521-85181 [30]. 978-0-521-85181-7 [46].

Accuracy [49]. Adapting [112]. advances [45]. after [108, 111]. Alexander [82, 84, 80, 90]. Analog [115]. analytical [48]. Ancient [30, 82, 35, 84, 41, 38, 80, 46, 47, 98, 90, 1, 27, 102, 103, 104, 114, 29, 70, 26, 107]. ancients [32]. Anomaly [39, 62]. Antikythera [52, 7, 8, 10, 82, 9, 15, 84, 114, 79, 6, 80, 81, 12, 13, 14, 90, 59, 63, 2, 48, 53, 21, 99, 100, 54, 65, 78, 25, 35, 83, 91, 92, 23, 49, 60, 67, 55, 109, 68, 36, 39, 93, 27, 69, 94, 95, 102, 103, 104, 56, 40, 57, 96, 97, 85, 37, 72, 73, 32, 76, 110, 11, 16, 51, 44, 45, 5, 86, 70, 87, 34, 88, 89, 98, 105, 108, 111, 58, 61, 62, 3, 4]. Antiquity [14]. Apokatastasis [111, 108]. Application [54]. Appreciation [16]. April [113]. Archaeological [25]. Archimedes [71]. Arithmetic [113]. Assembling [105]. Assessment [49]. astonishing [70]. Astronomical [59, 72, 110, 81, 27, 70, 17, 19, 20]. Astronomy [29, 23, 41, 30, 46, 38, 47]. Atom [35]. Atomic [85]. Austin [113]. Author [102, 103].

B.C. [10, 5, 3, 4, 8, 9, 6, 12, 13, 14]. Babylonian [93]. Back [63, 97, 96]. BC [7]. Before [25]. Bis [15]. Bomb [85]. Book [30, 8, 10, 82, 36, 11, 38, 6, 80, 81, 12, 13, 46, 47, 14, 90]. Building [71]. buried [77]. Byzantine [18].

ca [7, 8, 10, 9, 6, 12, 13, 5, 14, 3, 4]. Calculating [86]. calculator [27, 45]. Calendar [7, 8, 10, 9, 6, 12, 13, 14, 4, 35, 5, 3]. Calendars [30, 38, 46, 47, 31, 29, 17, 19, 20]. calendrical [18]. Calibration [108, 111]. Cambridge [30, 46]. Case [53, 109, 48]. centuries [45]. Century [36, 43]. Century-Long [36, 43]. charts [35]. Chronological [53]. Classical [30, 38, 46, 47, 29]. Classroom [112]. clock [31]. clockwork [33]. combines [31]. Complex [31]. Compound [61, 91]. Computation [50]. Computer [7, 8, 10, 9, 50, 36, 113, 6, 12, 13, 14, 34, 1, 4, 42, 43, 5, 33, 70, 3]. Computing [115, 22, 60, 106]. Conclusions [88]. Constantine [79]. Construction [63, 55, 98]. Correction [24, 94, 102, 103, 97]. Cosmos [82, 102, 103, 104, 84, 80, 90, 35, 114]. Cultural [109]. current [87].

D [7, 105]. dark [89]. Daryn [30, 46]. Date [78, 108, 111]. days [64]. Decode [50]. Decoding [27, 72, 73, 42, 43, 70, 81, 36]. Deepening [83]. Derek [8, 10, 9, 16, 6, 13, 14, 85, 12]. Desolla [16]. Determination [55]. determine [86]. Device [54, 72, 40, 73, 81]. Dial [75, 58, 73]. differences [56]. Digital [115]. Discover [36, 43]. display [73]. Displays [39]. Documentation [109]. Draconic [105]. driving [86].

Early [18, 60]. Eastern [30, 38, 46, 47, 29]. eclipse [65, 94, 95, 96, 97]. eds [79]. Efthymios [79]. Egyptians [19]. epicyclic [91]. epoch [65]. Euctemon [17]. Events [59].

Famed [68, 76]. fields [48]. Final [78]. Finding [85]. Finds [53]. First [25, 50, 110, 42]. found [76]. Fragment [105]. Front [75, 58, 73]. Functional [88]. gadget [35]. game [35]. Gear [49, 37, 61, 86]. Gear-Trains [61]. gearing [91, 18, 105]. Gears [5, 14, 3, 4, 55, 12, 7, 8, 10, 9, 11, 6, 13]. Geminos [20]. geometrical [55]. giant [77]. Greece [48, 53, 25, 26]. Greek [23, 27, 102, 103, 104, 70, 1, 17, 19, 20]. Greeks [7, 8, 9, 6, 12, 13, 14, 5, 3, 4, 10, 11].

Hanny [56]. hardback [46]. Heavens [36, 42, 43]. Heritage [109]. High [26]. Hipparchus [92, 40]. Historical [112]. History [14, 22, 60]. Holistic [109]. Hong [81]. Hong-Sen [81]. Hugh [101]. Human [76].

IEEE [113]. II [60]. Images [66]. Indications [61]. influence [92]. Initial [49, 108, 111]. Inscription [96, 97]. Inscriptions [75, 79, 44]. insights [56]. inspiration [41]. Intensive [53]. interactivity [83]. interior [62]. ISBN [30, 46]. IV [19].

Jian [81]. Jian-Liang [81]. John [101]. Jones [80, 90, 82, 84].

kept [35]. Keynote [60]. Knowledge [112, 87]. Known [50, 27].

landscape [48]. Lego [52]. Lehoux [30, 46]. Liang [81]. Like [85]. Lin [81]. location [86]. Long [36, 43]. lost [28, 32, 62]. lunar [40].

Machine [110, 98]. mathematical [18]. Measuring [53]. Mechanical [41, 111, 106, 60, 67, 108]. Mechanism [7, 8, 10, 54, 82, 9, 109, 102, 103, 104, 84, 85, 16, 6, 80, 12, 13, 14, 108, 111, 61, 90, 4, 65, 35, 67, 93, 94, 95, 73, 5, 70, 62, 3, 59, 63, 52, 21, 99, 100, 78, 83, 91, 92, 23, 49, 60, 55, 36, 39, 27, 69, 56, 40, 57, 96, 97, 114, 37, 11, 51, 44, 86, 87, 34, 88, 89, 98, 105, 58, 62, 79]. Mechanisms [72, 81]. Mediterranean [48]. Metonic [63]. Milestones [115]. Model [102, 103, 104, 55]. models [40]. modern [45]. month [66]. moon [73]. motion [62, 20]. motions [40]. Movable [109]. multimedia [74]. mystery [23, 42].

Near [30, 38, 46, 47, 29]. Near-Eastern [30, 47]. necessary [55]. Nicolaidis [79]. Nobel [77]. Notes [21].

Obituary [16]. Objects [109]. October [64, 66]. Old [36, 43]. Olympics [35]. Opening [85]. operation [105]. operational [55]. Our [87].

parameters [55]. parapegma [17, 19, 20, 59, 75]. Parapegmata [30, 38, 46, 47, 29, 47]. parts [98]. perpetual [19]. Perspective [34]. phase [73]. Phases [37]. picks [74]. Pin [54, 40]. Pin-and-Slot [54]. Planetary [39, 61]. Planets [54]. Plate [63, 96, 97]. Pointer [63]. Portable [82, 84, 90, 114, 80]. Pp [46, 14, 30]. prediction [94, 95]. predictor [65]. Press [30, 46]. Price [10, 12, 14, 7, 8, 9, 85, 16, 6, 13]. prizes [77].

Proceedings [113]. Production [112]. Ptolemy [40]. Publications [14]. Pyramid [85].

Rechner [15]. Reconstructing [69]. Reconstruction [88, 62]. Reflections [44]. Related [30, 38, 46, 47, 29]. Revealing [82, 84, 80, 90, 114]. Review [30, 7, 8, 10, 82, 36, 11, 38, 6, 80, 81, 12, 13, 46, 47, 14, 90, 9, 84]. Revising [94, 95]. revisited [96, 97]. role [105]. Romaka [20].

Saga [36]. Saros [108, 111]. scheme [94, 95, 96, 97]. Science [14, 34, 45]. Scientific [82, 84, 80, 90, 114]. Search [36, 28, 43]. secret [32]. Secrets [37, 43, 36]. Seiradakis [101]. Sen [81]. Seven [64]. shades [89]. shafts [86]. Shipwreck [2, 76]. Siddhānta [20]. similarities [56]. skeleton [76]. Skordoulis [79]. Slot [54, 40]. Societies [30, 29, 46, 38, 47]. Solar [39, 93, 62]. Solla [8, 10, 9, 85, 6, 12, 13, 14]. solving [42]. sphere [71]. Spiral [111, 108]. Spirals [63]. still [23]. structure [48]. Study [53, 48]. Sun [20]. Superior [54]. Survey [53, 25]. Symposium [113]. synthesis [62]. System [37].

tables [19]. Tale [33]. tech [26]. Technology [22]. telescope [77]. Terraced [48]. Texas [113]. Texts [30, 38, 46, 47, 29]. their [19]. theory [93]. time [35, 28]. tools [98]. Top [74]. torque [86]. Trains [49, 61]. treasure [77]. treasures [68].

Uncertainty [53]. Underwater [25]. universe [67]. University [30, 46]. Unraveling [37]. USA [113]. Using [50].

V [20]. Venus [91]. via [83]. voorwerp [56]. vs [40].

Weather [30, 38, 46, 47, 29]. wheelwork [18]. Wonder [82, 107, 84, 80, 90, 114]. World [38, 46, 47, 114, 29, 42, 30, 107, 82, 84, 80, 90]. wreck [68].

xiv [30, 46].

Yan [81]. Year [36, 43]. yields [68]. York [46, 14].

Z [15]. Zuse [15].

References

deSollaPrice:1959:AGC

[1] Derek J. de Solla Price. An ancient Greek computer. Scientific American, 200(6):60–67, June 1959. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v200/n6/pdf/scientificamerican0659-
60.pdf.

Anonymous:1965:AS

[2] Anonymous. The Antikythera shipwreck. Nature, 207(5004):1338, September 1, 1965. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

deSollaPrice:1974:GGA

[3] Derek John de Solla Price. Gears from the Greeks: the Antikythera mechanism, a calendar computer from ca. 80 B.C. Transactions of the American Philosophical Society, New Series, 64(7):1–70, November 1974. CODEN TAPSAY. ISBN 0-87169-647-9. ISSN 0065-9746. URL http://www.

jstor.org/stable/1006146.

deSollaPrice:1974:GGAa

[4] Derek de Solla Price. Gears from the Greeks. The Antikythera mechanism: a calendar computer from ca. 80 B.C. Transactions of the American Philosophical Society, New Series, 64(7):1–70, November 1974. CODEN TAP-

SAY. ISSN 0065-9746. URL http://www.jstor.org/stable/1006146.

Price:1974:GGAb

[5] Derek J. de Solla (Derek John de Solla) Price. Gears from the Greeks: the Antikythera mechanism, a calendar computer from ca. 80 B.C., volume 64(7) of Transactions of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1974. ISBN 0-87169-647-9. ISSN 0065-9746. 70 pp. LCCN QB107 .P74. US$5.00.

Nash:1975:BRB

[6] Robert A. Nash. Book review: Gears from the Greeks: The Antikythera Mechanism — A Calendar Computer from ca. 80 B.C. by Derek de Solla Price. American Scientist, 63(5):594, September 1975. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http:/

/www.jstor.org/stable/27845769.

Ashfaque:1976:RBG

[7] S. M. Ashfaque. Review: Gears from Greeks — Antikythera Mechanism — Calendar Computer from ca 80 BC — by D. D. Price. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 20(2):175–176, ???? 1976. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Ballin:1976:BRB

[8] Theodore N. Ballin. Book review: Gears from the Greeks, The Antikythera Mechanism: A Calendar Computer from ca. 80 B.C. by Derek de Solla Price. The Classical World, 70(3):202, November 1976. ISSN 0009-8418 (print), 1558-9234 (electronic). URL http://www.jstor.org/stable/

4348621.

Drachmann:1976:GGA

[9] A. G. Drachmann. Gears from the Greeks: The Antikythera Mechanism — a Calendar Computer from ca. 80 B.C. by Derek de Solla Price (review). Technology and Culture, 17(1):112–116, January 1976. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL http://www.jstor.org/stable/3103259;   https://muse.jhu.
edu/pub/1/article/891657/pdf.

Brookes:1977:BRB

[10] A. M. P. Brookes. Book review: Gears from the Greeks, by Derek de Solla Price. The Antikythera Mechanism: A Calendar Computer from ca. 80 B.C. The Classical Review, 27(1):94–95, ???? 1977. CODEN ???? ISSN 0009-840X (print), 1464-3561 (electronic). URL http://www.jstor.org/

stable/710979.

Mercier:1977:BRA

[11] Raymond Mercier. Book review: The Antikythera Mechanism, Gears from the Greeks. Journal for the History of Astronomy, 8(2):143–145, June 1977. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

North:1977:BRD

[12] J. D. North. Book review: Derek de Solla Price: Gears from the Greeks: The Antikythera Mechanism. A Calendar Computer from ca. 80 B.C. Isis, 68(1):142–143, March 1977. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/230405.

Oleson:1977:BRB

[13] John Peter Oleson. Book review: Gears from the Greeks: The Antikythera Mechanism — a Calendar Computer from ca. 80 B.C. by Derek de Solla Price. Phoenix, 31(3):271–273, Autumn 1977. ISSN 0031-8299 (print),

1929-4883 (electronic). URL http://www.jstor.org/stable/1087110.

Smith:1978:BRA

[14] Norman A. F. Smith. Book review: Antiquity Gears from the Greeks: The Antikythera Mechanism — a Calendar Computer from ca. 80 B.C. by Derek de Solla Price. New York: Science History Publications, 1975. Pp.

70. $8.50. British Journal for the History of Science, 11(1):77–78, March 1978. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/stable/4025619.

Handler:1979:RBZ

[15] Wolfgang Händler. Rechner Von A Bis Z — Von Antikythera Bis Zuse. In K. H. Böhling and P. P. Spies, editors, GI – 9. Jahrestagung, pages 1–15. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979.

Morris:1984:PDD

[16] L. R. Morris. Price, Derek, Desolla and the Antikythera mechanism — an appreciation 1922–1983 — obituary. IEEE Micro, 4(1):15–21, January/ February 1984. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

vanderWaerden:1984:GACa

[17] B. L. van der Waerden. Greek astronomical calendars I. The parapegma of Euctemon. Archive for History of Exact Sciences, 29(2):101– 114, June 1984. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=29&issue=2&spage=101.

Maddison:1985:EMW

[18] Francis Maddison. Early mathematical wheelwork: Byzantine calendrical gearing. Nature, 314(6009):316–317, March 28, 1985. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

vanderWaerden:1985:GACa

[19] B. L. van der Waerden. Greek astronomical calendars IV. The parapegma of the Egyptians and their “perpetual tables”. Archive for History of Exact Sciences, 32(2):95–104, June 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0003-9519&volume=32&issue=2&spage=
95.

vanderWaerden:1985:GACb

[20] B. L. van der Waerden. Greek astronomical calendars V. The motion of the Sun in the parapegma of Geminos and in the Romaka– Siddhānta. Archive for History of Exact Sciences, 34(3):231–239, September 1985. CODEN AHESAN. ISSN 0003-9519 (print), 1432-

0657 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0003-9519&volume=34&issue=3&spage=231.

Bromley:1986:NAM

[21] Allan G. Bromley. Notes on the Antikythera Mechanism. Centaurus: An International Journal of the History of Science and its Cultural Aspects, 29(1):5–27, March 1986. CODEN CENTA4. ISSN 0008-8994 (print), 1600-0498 (electronic).

Williams:1997:HCT

[22] Michael R. Williams. A History of Computing Technology. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, second edition, 1997. ISBN 0-8186-7739-2 (hardcover). xi + 426 pp. LCCN QA76.17 .W55 1997. URL http://catalogue.bnf.fr/

ark:/12148/cb37514972q.

Edmunds:2000:AMS

[23] M. G. Edmunds and P. Morgan. The Antikythera Mechanism: still a mystery of Greek astronomy? Astronomy and Geophysics, 41(6):6.10– 6.17, December 2000. CODEN ASGEF5. ISSN 1366-8781 (print), 1468-4004 (electronic).

Anonymous:2006:C

[24] Anonymous. Correction. Nature, 444(7120):699, December 6, 2006. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). See [26].

Catsambis:2006:BAF

[25] Alexis Catsambis. Before Antikythera: the first underwater archaeological survey in Greece. International Journal of Nautical Archaeology, 35(1): 104–107, April 2006. ISSN 1057-2414 (print), 1095-9270 (electronic).

Charette:2006:HTA

[26] François Charette. High tech from Ancient Greece. Nature, 444(7119): 551–552, November 1, 2006. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). See correction [24].

Freeth:2006:DAG

[27] T. Freeth, Y. Bitsakis, X. Moussas, J. H. Seiradakis, A. Tselikas, H. Mangou, M. Zafeiropoulou, R. Hadland, D. Bate, A. Ramsey, M. Allen, A. Crawley, P. Hockley, T. Malzbender, D. Gelb, W. Ambrisco, and M. G. Edmunds. Decoding the ancient Greek astronomical calculator known as the Antikythera Mechanism. Nature, 444(7119):587–591, November 1, 2006. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Marchant:2006:SLT

[28] Jo Marchant. In search of lost time. Nature, 444(7119):534–538, November 1, 2006. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Lehoux:2007:AWC

[29] Daryn Lehoux. Astronomy, weather, and calendars in the ancient world: parapegmata and related texts in classical and Near Eastern societies. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-85181-5 (hardcover). xiv + 566 pp. LCCN QC999

.L44 2007. URL http://www.loc.gov/catdir/enhancements/fy0703/
2006038547-d.html;   http://www.loc.gov/catdir/enhancements/
fy0703/2006038547-t.html;   http://www.loc.gov/catdir/enhancements/
fy0729/2006038547-b.html.

Acerbi:2008:BRD

[30] Fabio Acerbi. Book review: Daryn Lehoux, Astronomy, Weather, and Calendars in the Ancient World. Parapegmata and Related Texts in Classical and Near-Eastern Societies. Cambridge: Cambridge University Press, 2007. xiv + 566 pp., ISBN 978-0-521-85181. Nuncius, 23(1):134–135, ???? 2008. CODEN ???? ISSN 0394-7394 (print), 1825-3911 (electronic).

URL http://booksandjournals.brillonline.com/content/10.1163/
182539108x00111.

Ball:2008:CCC

[31] Philip Ball. Complex clock combines calendars. Nature, 454(7204):561, July 30, 2008. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Marchant:2008:ALS

[32] Jo Marchant. The Antikythera: lost secret of the ancients. New Scientist, 200(2686):36–40, 2008. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL https://www.sciencedirect.com/

science/article/pii/S0262407908631619.

Robinson:2008:TCC

[33] Andrew Robinson. Tale of a clockwork computer. Nature, 455(7215): 867–868, October 15, 2008. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Spinellis:2008:AMC

[34] Diomidis Spinellis. The Antikythera Mechanism: a computer science perspective. Computer, 41(5):22–27, May 2008. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cowen:2009:ACA

[35] Ron Cowen. Atom & cosmos: Ancient gadget charts game time: Antikythera mechanism kept a calendar of the Olympics. Science News (Washington, DC), 174(5):10, September 2009. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic).

Evans:2009:BRS

[36] James Evans. Book review: The saga of the Antikythera Mechanism, Decoding the Heavens: a 2000-Year-Old Computer and the Century-Long Search to Discover its Secrets. Journal for the History of Astronomy, 40 (3):362–364, August 2009. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Koetsier:2009:PUS

[37] Teun Koetsier. Phases in the unraveling of the secrets of the gear system of the Antikythera Mechanism. In Hong-Sen Yan and Marco Ceccarelli, editors, International Symposium on History of Machines and Mechanisms, pages 269–294. Springer Netherlands, Dordrecht, The Netherlands, 2009.

Montelle:2009:BRB

[38] Clemency Montelle. Book review: Astronomy, Weather, and Calendars in the Ancient World: Parapegmata and Related Texts in Classical and Near Eastern Societies. Isis, 100(4):896–897, December 2009. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/10.1086/652047.

Evans:2010:SAP

[39] James Evans, Christián C. Carman, and Alan S. Thorndike. Solar anomaly and planetary displays in the Antikythera Mechanism. Journal for the History of Astronomy, 41(1):1–39, February 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Gourtsoyannis:2010:HVP

[40] Elias Gourtsoyannis. Hipparchus vs. Ptolemy and the Antikythera Mechanism: Pin slot device models lunar motions. Advances in Space Research, 46(4):540–544, 2010. CODEN ASRSDW. ISSN 0273-1177 (print), 1879-

1948 (electronic). URL https://www.sciencedirect.com/science/

article/pii/S0273117709006036. Advances in Space Environment Research.

Marchant:2010:AAM

[41] Jo Marchant. Ancient astronomy: Mechanical inspiration. Nature, 468 (7323):496–498, November 24, 2010. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Marchant:2010:DHS

[42] Josephine Marchant. Decoding the heavens: solving the mystery of the world’s first computer. William Heinemann, London, UK, 2010. ISBN 0-434-01835-X (hardcover). 328 + 8 pp. LCCN QB107 .M37

2010. URL http://site.ebrary.com/id/10372916;   http://www.
decodingtheheavens.com/reviews.aspx.

Marchant:2010:DHY

[43] Josephine Marchant. Decoding the heavens: a 2,000-year-old-computer — and the century-long search to discover its secrets. Da Capo Press, Cambridge, MA, USA, 2010. ISBN 0-306-81861-2 (paperback). 328 + 8

pp. LCCN QB107 .M38 2010. URL http://www.decodingtheheavens.
com/reviews.aspx.

Papathanassiou:2010:RAM

[44] Maria K. Papathanassiou. Reflections on the Antikythera Mechanism inscriptions. Advances in Space Research, 46(4):545–551, 2010. CODEN ASRSDW. ISSN 0273-1177 (print), 1879-1948 (electronic). URL https:/

/www.sciencedirect.com/science/article/pii/S027311770900684X.

Advances in Space Environment Research.

Pastore:2010:ACA

[45] Giovanni Pastore. Antikythera Calculator advances modern science of 19 centuries. Advances in Space Research, 46(4):552–556, 2010. CODEN ASRSDW. ISSN 0273-1177 (print), 1879-1948 (electronic). URL https:/

/www.sciencedirect.com/science/article/pii/S027311771000236X.

Advances in Space Environment Research.

Robson:2010:BRD

[46] Eleanor Robson. Book review: Daryn Lehoux, Astronomy, Weather, and Calendars in the Ancient World: Parapegmata and Related Texts in Classical and Ancient Near Eastern Societies. Cambridge and New York:

Cambridge University Press, 2007. Pp. xiv + 566. ISBN 978-0-521-85181-7. £65.00 (hardback). British Journal for the History of Science, 43(2): 288–289, June 2010. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Sidoli:2010:BRP

[47] Nathan Sidoli. Book review: Parapegmata: Astronomy, Weather, and Calendars in the Ancient World: Parapegmata and Related Texts in Classical and Near-Eastern Societies. Journal for the History of Astronomy, 41(1):124–126, February 2010. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Bevan:2011:TFM

[48] Andrew Bevan and James Conolly. Terraced fields and Mediterranean landscape structure: an analytical case study from Antikythera, Greece. Ecological Modelling, 222(7):1303–1314, April 10, 2011. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http:/

/www.sciencedirect.com/science/article/pii/S0304380010006824.

Edmunds:2011:IAA

[49] M. G. Edmunds. An initial assessment of the accuracy of the gear trains in the Antikythera Mechanism. Journal for the History of Astronomy, 42 (3):307–320, August 2011. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Edmunds:2011:UCD

[50] Michael Edmunds and Tony Freeth. Using computation to decode the first known computer. Computer, 44(7):32–39, July 2011. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moussas:2011:AM

[51] Xenophon Moussas. The Antikythera Mechanism. In Kokkotas et al. [112], pages 113–128. ISBN 94-6091-349-0. LCCN Q126.9 .A33 2011;

Q181.A1 .K384 2010. URL http://www.springerlink.com/content/
978-94-6091-349-5.

Anonymous:2012:LAM

[52] Anonymous. Lego Antikythera Mechanism. Nature, December 2012. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Bevan:2012:MCU

[53] A. Bevan, J. Conolly, C. Hennig, A. Johnston, A. Quercia, L. Spencer, and J. Vroom. Measuring chronological uncertainty in intensive survey finds: a case study from Antikythera, Greece. Archaeometry, 55(2):312– 328, March 2012. CODEN ARCHAG. ISSN 0003-813X (print), 1475-4754 (electronic).

Carman:2012:PSD

[54] Christián C. Carman, Alan Thorndike, and James Evans. On the pinand-slot device of the Antikythera mechanism, with a new application to the superior planets. Journal for the History of Astronomy, 43(1):93–116, February 2012. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Efstathiou:2012:DGG

[55] K. Efstathiou, A. Basiakoulis, M. Efstathiou, M. Anastasiou, and J. H. Seiradakis. Determination of the gears geometrical parameters necessary for the construction of an operational model of the Antikythera Mechanism. Mechanism and Machine Theory, 52:219–231, 2012. CODEN MHMTAS. ISSN 0094-114X (print), 1873-3999 (electronic). URL https:/

/www.sciencedirect.com/science/article/pii/S0094114X12000316.

Garrett:2012:HVA

[56] Michael A. Garrett. Hanny’s voorwerp and the Antikythera Mechanism — similarities, differences and insights. arxiv.org, ??(??):1–8, November

23, 2012. URL https://arxiv.org/abs/1211.5487.

Hannah:2012:AM

[57] Robert Hannah. Antikythera Mechanism, October 2012.

Wright:2012:FDA

[58] Michael T. Wright. The front dial of the Antikythera Mechanism. In Teun Koetsier and Marco Ceccarelli, editors, History of Mechanism and Machine Science, pages 279–292. Springer Netherlands, Dordrecht, The Netherlands, 2012.

Anastasiou:2013:AEP

[59] Magdalini Anastasiou, John H. Seiradakis, James Evans, Stella Drougou, and Kyriakos Efstathiou. The astronomical events of the Parapegma of the Antikythera Mechanism. Journal for the History of Astronomy, 44(2): 173–A10, May 2013. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Edmunds:2013:AME

[60] Michael G. Edmunds. Keynote II: The Antikythera Mechanism and the early history of mechanical computing. In IEEE [113], page 79. ISBN 0-7695-4957-8. ISSN 1063-6889. LCCN QA76.9.C62 S95 2013.

Wright:2013:AMC

[61] M. T. Wright. The Antikythera mechanism: Compound gear-trains for planetary indications. Almagest: International Journal for the History of Scientific Ideas, 4(2):4–31, 2013. CODEN ???? ISSN 1792-

2593. URL https://www.brepolsonline.net/doi/epdf/10.1484/J.
ALMAGEST.1.103717.

Yan:2013:RSL

[62] Hong-Sen Yan and Jian-Liang Lin. Reconstruction synthesis of the lost interior mechanism for the solar anomaly motion of the Antikythera Mechanism. Mechanism and Machine Theory, 70:354–371, 2013. CODEN MHM-TAS. ISSN 0094-114X (print), 1873-3999 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S0094114X13001638.

Anastasiou:2014:AMC

[63] M. Anastasiou, J. H. Seiradakis, C. C. Carman, and K. Efstathiou. The Antikythera Mechanism: The construction of the metonic pointer and the back plate spirals. Journal for the History of Astronomy, 45(4):418–441, November 2014. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Anonymous:2014:SDO

[64] Anonymous. Seven days: 10–16 October 2014. Nature, 514(7522):278–279, October 15, 2014. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Carman:2014:EAM

[65] Christián C. Carman and James Evans. On the epoch of the Antikythera mechanism and its eclipse predictor. Archive for History of Exact Sciences, 68(6):693–774, November 2014. CODEN AHE-SAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://

link.springer.com/article/10.1007/s00407-014-0145-5;   http:
//www.scientificcomputing.com/news/2014/12/worlds-oldest-computer-
ancient-greek-antikythera-mechanism-100-years-older-previously-
believed.

Cressey:2014:IMO

[66] Daniel Cressey and Davide Castelvecchi. Images of the month: October 2014. Nature, October 2014. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Edmunds:2014:AMM

[67] M. G. Edmunds. The Antikythera mechanism and the mechanical universe. Contemporary Physics, 55(4):263–285, 2014. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Ehrenberg:2014:FAW

[68] Rachel Ehrenberg. Famed Antikythera wreck yields more treasures. Nature, October 10, 2014. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Freeth:2014:RAM

[69] Tony Freeth. Reconstructing the Antikythera Mechanism. In Clive L. N. Ruggles, editor, Handbook of Archaeoastronomy and Ethnoastronomy, pages 1603–1624. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., July 2014.

Seiradakis:2014:AMD

[70] John H. Seiradakis. The Antikythera mechanism: decoding an astonishing ancient Greek astronomical computer. Physica Medica, 30:e2, 2014. ISSN 1120-1797 (print), 1724-191X (electronic). URL https://

www.sciencedirect.com/science/article/pii/S1120179714001793.

Baker:2015:BSA

[71] Noah Baker. Building the sphere of Archimedes. Nature, September 2015. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Lin:2015:DMAb

[72] Jian-Liang Lin and Hong-Sen Yan. Decoding the Mechanisms of Antikythera Astronomical Device. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-662-48445-5 (hardcover), 3-662-48447-1 (e-book). x + 281 + 172 + 141 pp. LCCN QB107 .L56 2016; QC5.53.

Lin:2015:DMP

[73] Jian Liang Lin and Hong Sen Yan. Decoding the moon phase display device over the front dial of the Antikythera mechanism. Chinese Journal of Mechanical Engineering (English Edition), 28(4):676–683, July 2015. CODEN CJMEER. ISSN 1000-9345 (print), 2192-8258 (electronic).

Anonymous:2016:TMP

[74] Anonymous. Top multimedia picks of 2016. Nature, December 2016. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Bitsakis:2016:FDP

[75] Y. Bitsakis and A. Jones. The front dial and Parapegma inscriptions. Almagest: International Journal for the History of Scientific Ideas, 7(1): 68–137, 2016. CODEN ???? ISSN 1792-2593. URL https://www.

brepolsonline.net/doi/epdf/10.1484/J.ALMAGEST.5.110734.

Marchant:2016:HSF

[76] Jo Marchant. Human skeleton found on famed Antikythera shipwreck. Nature, 537(7621):462–463, September 19, 2016. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Anonymous:2017:NPG

[77] Anonymous. Nobel prizes, giant telescope and buried treasure. Nature, 550(7674):12–13, October 5, 2017. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Carman:2017:FDA

[78] Christián C. Carman. The final date of the Antikythera Mechanism. Journal for the History of Astronomy, 48(3):312–323, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic). URL http://

journals.sagepub.com/doi/full/10.1177/0021828617721553.

Koetsier:2017:ENC

[79] Teun Koetsier. Efthymios Nicolaidis; Constantine Skordoulis, eds. The Inscriptions of the Antikythera Mechanism. Isis, 108(3):685–686, September 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Nicolaidis:2017:BRA

[80] Efthymios Nicolaidis. Book review: Alexander Jones, A portable Cosmos. Revealing the Antikythera Mechanism, Scientific Wonder of the Ancient World. Almagest: International Journal for the History of Scientific Ideas, 8(1):133–137, 2017. CODEN ???? ISSN 1792-2593. URL https://www.

brepolsonline.net/doi/epdf/10.1484/J.ALMAGEST.5.113701.

Nicolaidis:2017:BRJ

[81] Efthymios Nicolaidis. Book review: Jian-Liang Lin; Hong-Sen Yan. Decoding the Mechanisms of Antikythera Astronomical Device. Isis, 108(1): 172–173, March 2017. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Carman:2018:BRP

[82] Christián C. Carman. Book review: A Portable Cosmos: Revealing the Antikythera Mechanism, Scientific Wonder of the Ancient World, by Alexander Jones. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 33(3):189–190, 2018. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/17498430.2018.1457342.

Diolatzis:2018:DAM

[83] Ioannis S. Diolatzis and Gerasimos Pavlogeorgatos. Deepening to Antikythera Mechanism via its interactivity. Digital Applications in Archaeology and Cultural Heritage, 8:10–26, 2018. ISSN 2212-0548. URL https:/

/www.sciencedirect.com/science/article/pii/S221205481730022X.

Henderson:2018:PCR

[84] Georgina Henderson. A Portable Cosmos: Revealing the Antikythera Mechanism, Scientific Wonder of the Ancient World by Alexander Jones (review). Technology and Culture, 59(3):789–790, July 2018. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/703149.

Jones:2018:LOP

[85] Alexander Jones. “Like opening a pyramid and finding an atomic bomb”: Derek de Solla Price and the Antikythera mechanism. Proceedings of the American Philosophical Society, 162(3):259–294, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2019-03/attachments/Jones.pdf.

Roumeliotis:2018:CTS

[86] Manos Roumeliotis. Calculating the torque on the shafts of the Antikythera Mechanism to determine the location of the driving gear. Mechanism and Machine Theory, 122:148–159, 2018. CODEN MHMTAS. ISSN 0094-114X (print), 1873-3999 (electronic). URL https://www.sciencedirect.

com/science/article/pii/S0094114X17315586.

Seiradakis:2018:OCK

[87] J. H. Seiradakis and M. G. Edmunds. Our current knowledge of the Antikythera Mechanism. Nature Astronomy, 2(1):35–42, January 2018.

Voulgaris:2018:CFR

[88] Aristeidis Voulgaris, Christophoros Mouratidis, and Andreas Vossinakis. Conclusions from the functional reconstruction of the Antikythera Mechanism. Journal for the History of Astronomy, 49(2):216–238, May 2018. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556

(electronic). URL http://journals.sagepub.com/doi/full/10.1177/
0021828618762460.

Voulgaris:2018:DSA

[89] Aristeidis Voulgaris, Christophoros Mouratidis, and Andreas Vossinakis. The dark shades of the Antikythera Mechanism. Journal of Radioanalytical and Nuclear Chemistry, 318(3):1881–1891, October 2018. CODEN JRNCDM. ISSN 0236-5731 (print), 1588-2780 (electronic).

Yan:2018:BRA

[90] Hong-Sen Yan. Book review: Alexander Jones. A Portable Cosmos: Revealing the Antikythera Mechanism, Scientific Wonder of the Ancient World. Isis, 109(4):827–828, December 2018. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Diolatzis:2019:AMC

[91] Ioannis S. Diolatzis and Gerasimos Pavlogeorgatos. Antikythera Mechanism: a compound epicyclic gearing for Venus. Digital Applications in Archaeology and Cultural Heritage, 12:e00089, 2019. ISSN 2212-

0548. URL https://www.sciencedirect.com/science/article/pii/
S2212054818300341.

Diolatzis:2019:IHA

[92] Ioannis S. Diolatzis and Gerasimos Pavlogeorgatos. The influence of Hipparchus in Antikythera Mechanism. New Astronomy, 67:29–39, 2019. CODEN NEASFS. ISSN 1384-1076 (print), 1384-1092 (elec-

tronic). URL https://www.sciencedirect.com/science/article/
pii/S1384107618302227.

Evans:2019:BST

[93] James Evans and Christián C. Carman. Babylonian solar theory on the Antikythera mechanism. Archive for History of Exact Sciences, 73(6): 619–659, November 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

Freeth:2019:CRE

[94] Tony Freeth. Correction: Revising the eclipse prediction scheme in the Antikythera mechanism. Palgrave Communications, 5(1), March 2019.

Freeth:2019:REP

[95] Tony Freeth. Revising the eclipse prediction scheme in the Antikythera mechanism. Palgrave Communications, 5(1), January 2019.

Iversen:2019:BPI

[96] Paul Iversen and Alexander Jones. The Back Plate Inscription and eclipse scheme of the Antikythera Mechanism revisited. Archive for History of Exact Sciences, 73(5):469–511, September 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s00407-019-00229-9.pdf. See correction

[97].

Iversen:2019:CBP

[97] Paul Iversen and Alexander Jones. Correction to: The Back Plate Inscription and eclipse scheme of the Antikythera Mechanism revisited. Archive for History of Exact Sciences, 73(5):513–516, September 2019. CODEN AHESAN. ISSN 0003-9519 (print), 1432-0657 (electronic).

URL http://link.springer.com/content/pdf/10.1007/s00407-019-
00234-y.pdf. See [96].

Voulgaris:2019:AMT

[98] Aristeidis Voulgaris, Christophoros Mouratidis, and Andreas Vossinakis. Ancient machine tools for the construction of the Antikythera Mechanism parts. Digital Applications in Archaeology and Cultural Heritage, 13:e00092, 2019. ISSN 2212-0548. URL https://www.sciencedirect.

com/science/article/pii/S2212054818300353.

Bruderer:2020:AMa

[99] Herbert Bruderer. The Antikythera Mechanism. Communications of the ACM, 63(4):108–115, April 2020. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Bruderer:2020:AMb

[100] Herbert Bruderer. The Antikythera Mechanism. In Milestones in Analog and Digital Computing [115], pages 409–426. ISBN 3-030-40973-2, 3-030-40974-0 (e-book), 3-030-40975-9. LCCN QA76.17 .B78413 2020. URL

http://www.springerlink.com/content/978-3-030-40974-6.

Kalogera:2020:JHS

[101] Vassiliki Kalogera and Michael Kramer. John Hugh Seiradakis. Nature Astronomy, 4(7):639–640, June 2020.

Freeth:2021:ACMa

[102] Tony Freeth, David Higgon, Aris Dacanalis, Lindsay MacDonald, Myrto Georgakopoulou, and Adam Wojcik. Author correction: A model of the cosmos in the ancient Greek Antikythera mechanism. Scientific Reports, 11(1), March 2021. CODEN SRCEC3. ISSN 2045-2322. See [104].

Freeth:2021:ACMb

[103] Tony Freeth, David Higgon, Aris Dacanalis, Lindsay MacDonald, Myrto Georgakopoulou, and Adam Wojcik. Author correction: A model of the cosmos in the ancient Greek Antikythera mechanism. Scientific Reports, 11(1):??, August 2021. CODEN SRCEC3. ISSN 2045-2322. See [104]. From the introduction: “This original version of this article contained repeated errors, where the Greek character Υ (Upsilon) was incorrectly given as Ψ (Psi) and fractions were incorrectly shown as superscripts.”.

Freeth:2021:MCA

[104] Tony Freeth, David Higgon, Aris Dacanalis, Lindsay MacDonald, Myrto Georgakopoulou, and Adam Wojcik. A model of the cosmos in the ancient Greek Antikythera mechanism. Scientific Reports, 11(1), March 2021. CODEN SRCEC3. ISSN 2045-2322. URL https://www.nature.com/

articles/s41598-021-96382-9. See corrections [102, 103].

Voulgaris:2021:DGA

[105] Aristeidis Voulgaris, Christophoros Mouratidis, Andreas Vossinakis, and G. Bokovos. The Draconic gearing of the Antikythera Mechanism: Assembling the Fragment D, its role and operation. arxiv.org, ??(??):1–??,

March 24, 2021. URL https://arxiv.org/abs/2104.06181.

Yasuda:2021:MC

[106] Hiromi Yasuda, Philip R. Buskohl, Andrew Gillman, Todd D. Murphey, Susan Stepney, Richard A. Vaia, and Jordan R. Raney. Mechanical computing. Nature, 598(7879):39–48, October 6, 2021. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Freeth:2022:WAW

[107] Tony Freeth. Wonder of the Ancient World. Scientific American, 326 (1):24–33, January 2022. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Voulgaris:2022:ICD

[108] Aristeidis Voulgaris, Christophoros Mouratidis, and Andreas Vossinakis. The initial calibration date of the Antikythera mechanism after the Saros spiral mechanical Apokatastasis. arxiv.org, ??(??):1–29, March 28, 2022.

URL https://arxiv.org/abs/2203.15045.

Efstathiou:2023:HDM

[109] Kyriakos Efstathiou, Marianna Efstathiou, Alexandros Basiakoulis, Harriet Cliffen, Nenad Joncic, and Rafaella Georgiou. The holistic documentation of movable cultural heritage objects — the case of the Antikythera mechanism. In Marinos Ioannides and Petros Patias, editors, 3D Research Challenges in Cultural Heritage III: Complexity and Quality in Digitisation, volume 13125 of Lecture Notes in Computer Science, pages 103–130. Springer International Publishing, ????, 2023. ISBN 3-031-35593-8. ISSN 1611-3349.

Marini:2023:FAM

[110] Daniele L. R. Marini. The First Astronomical Machine: Antikythera, pages 161–177. Springer Nature, Cham, Switzerland, 2023. ISBN 3-303-13094-3-X, 3-031-30944-8 (e-book). ISSN 1614-659X (print), 2197-6651 (electronic). LCCN QB15 .M37 2023; QC6.9-9.

Voulgaris:2023:ICD

[111] Aristeidis Voulgaris, Christophoros Mouratidis, and Andreas Vossinakis. The initial calibration date of the Antikythera mechanism after the Saros spiral mechanical apokatastasis. Almagest: International Journal for the History of Scientific Ideas, 14(1):4–39, January 2023. CODEN ???? ISSN

1792-2593. URL https://www.brepolsonline.net/doi/epdf/10.1484/
J.ALMAGEST.5.134609.

Kokkotas:2011:AHK

[112] Panagiotis V. Kokkotas, Katerina S. Malamitsa, and Aikaterini A. Rizaki, editors. Adapting Historical Knowledge Production to the Classroom. SensePublishers, Rotterdam, The Netherlands, 2011. ISBN 94-6091-349-0. xii + 258 pp. LCCN Q126.9 .A33 2011; Q181.A1 .K384 2010.

IEEE:2013:PIS

[113] IEEE, editor. Proceedings of the 21st IEEE Symposium on Computer Arithmetic, Austin, Texas, USA, 8–10 April 2013. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN 0-7695-4957-8. ISSN 1063-6889. LCCN QA76.9.C62 S95 2013.

Jones:2017:PCR

[114] Alexander Jones. A portable cosmos: revealing the Antikythera Mechanism, scientific wonder of the ancient world. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2017. ISBN 0-19-973934-X. xiv + 288 + 8 pp. LCCN QB107 .J65 2017.

Bruderer:2020:MAD

[115] Herbert Bruderer. Milestones in Analog and Digital Computing. Springer International Publishing, Cham, Switzerland, third edition, 2020. ISBN 3-030-40973-2, 3-030-40974-0 (e-book), 3-030-40975-9. xxx + 2053 + 715 + 626 pp. LCCN QA76.17 .B78413 2020. URL http://www.springerlink.

com/content/978-3-030-40974-6.